OpenStudy (anonymous):

Find the inverse of f(x)= –9/x+8.

OpenStudy (anonymous):

Do you want the answer or a step by step solution?

OpenStudy (ranga):

y = -9/x + 8 To find inverse, switch x and y. That is, replace x by y and y by x. Express y as a function of x and that will be your inverse.

OpenStudy (anonymous):

step by step solution

OpenStudy (anonymous):

First set f(x) to Y. You now have: y=-9/x+8 do you agree?

OpenStudy (anonymous):

is it \[\frac{ -9 }{ x } + 8\] or \[\frac{ -9 }{ x+8 }\]

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

\[\it is \frac{ -9 }{ x}+8\]

OpenStudy (anonymous):

okay. f(x) is the same as Y. so now that we have y=-9/x+8, switch the x and y values.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So do it

OpenStudy (anonymous):

y=-9/y+8 ?

OpenStudy (anonymous):

close! you have to change that y in the beginning, too. so: x=-9/y+8 now, try to solve for x. the answer is f^-1(x) = -9/x-8 . I have to go, sorry